I need help on a question on my summer math worksheet for my upcoming year in 8th grade. The question says "Give the measure of the angle that is complementary to an angle that measures 71 degrees." all I ask is how to work it out. Thank you for your help.
Answer by jorel1380(3719)

m+n=90 degrees
m+71=90
m=90-71=19 degrees..